a debt of $2000 due in one year is to be repaid by a payment due two years from now and a final payment of $1000 three years from now. If the interest is at the rate of 4% compounded annually, then the payment due in two years is

the answer is $1118.46 but I need help on working it out
Answer #1

The equation for this cash flow in present value terms is given by

2000 (P/F, 4%, 1) = X (P/F, 4%, 2) + 1000 x (P/F, 4%, 3)

2000*(1.04)^-1 = X * (1.04)^-2 + 1000*(1.04)^-3

X = (2000*(1.04)^-1 - 1000*(1.04)^-3)/(1.04)^-2

= 1118.46

Hence payment due in two years is 1118.46